A rollercoaster is designed with a 10 m tall vertical circular loop. Calculate the minimum speed necessary to get around the loop without falling down.


1
Expert's answer
2021-11-09T16:36:18-0500
"F_g=F_c,""mg=\dfrac{mv^2}{r},""v_{min}=\sqrt{gr}=\sqrt{9.8\ \dfrac{m}{s^2}\times10\ m}=9.9\ \dfrac{m}{s}."